OpenStudy (anonymous):

I have a question, I will type it out using the equation feature, so wait a sec when you look at this :)

OpenStudy (anonymous):

sorry im lagging...almost done

OpenStudy (anonymous):

just a sec, ill try again, the bloody thing wont load

OpenStudy (anonymous):

\[\sqrt{18x}/4 +\sqrt{2x/4}\] No division, theyre fractions

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

add radicals

OpenStudy (anonymous):

3 root of x

OpenStudy (anonymous):

\[3x^{1/2}\]

OpenStudy (diyadiya):

sqrt 18x =3 sqrt2x \[\frac{3\sqrt{2x}}{2} + \frac{\sqrt{2x}}{2}\] \[=\frac{4\sqrt{2x}}{2}\] \[=2\sqrt{2x}\]

OpenStudy (anonymous):

so is that the final answer? the bottom

OpenStudy (anonymous):

(root of 18 * root of 2) * root of x root of 36 * root of x (6*root of x)/2 3 root of x

OpenStudy (anonymous):

Oh yes, I made a mistake with multiplication

OpenStudy (anonymous):

Sorry, regret my answer

OpenStudy (anonymous):

ok thanks Diya :) I will look over this and try more problems out ^_^

OpenStudy (anonymous):

are you sure? it said it was wrong

OpenStudy (anonymous):

\[2\sqrt{2x}\]

OpenStudy (anonymous):

no the top was \[5\sqrt{2x}\]

OpenStudy (anonymous):

the math thing checked it, then gave me a new problem >.<

OpenStudy (phi):

If the problem is \[ \frac{\sqrt{18x}}{4} +\sqrt{\frac{2x} {4}}\] then you would get \[ \frac{3\sqrt{2}}{4}+\frac{\sqrt{2x}}{2} \] \[ = \frac{3\sqrt{2}}{4}+\frac{2\sqrt{2x}}{4}=\frac{5\sqrt{2x}}{4} \]